Need help with the Excel PV function. The current price of the bond is 914.95 and the YTM are provided. The correct answers are on the right box but I dont get those numbers.

Face value Annual coupon rate Years to maturity $1,000 6.225% 526 27 28 29 30 Compounding Interval /Period 31 Annual 32 Semi-Annual 33 Quarterly 34 Monthly Compounding Interval Period $914.95 YTM 5.0% 8.34% 9.0% 854.25 S654.55 $622.16 1,053.61 $914.95$890.21 1,174.06 $1,091.89 $1,076.47 1,263.51 $1,232.76 $1,226.79 2 -$1,053.61- $914.95 -$890 12
Answer #1
A B C D E F G H I
2
3 Face Value $1,000
4 Annual Coupon rate 6.225%
5 Years to maturity 5
6 YTM 8.34%
7
8 Compounding Semi-annual
9
10 Compunding factor 2 (For Semi annual)
11
12 Periodic coupon $31.13 =D3*D4/D10
13 Number of period 10 =D10*D5
14 Periodic interest rate 4.17% =D6/D10
15
16 Price of Bond $914.95 =-PV(D14,D13,D12,D3)
17
18 Data table function of Excel can be used to find the values:
19 YTM
20 Compounding interval / Period $914.95 5.00% 8.34% 9.00%
21 Annual 1 1053.036 916.3055 892.0622
22 Semi-Annual 2 1053.606 914.9486 890.211
23 Quarterly 4 1053.898 914.2468 889.2517
24 Monthly 12 1054.095 913.7699 888.5991
25

Data Table

Data Table Row input cell: SDS6 Column input cell: SDS10 Cancel